# Rethinking Feature Backbone Fine-tuning for Remote Sensing Object Detection

### v1

@@ Abstract
Recently, numerous methods have achieved impressive performance in remote sensing object detection, relying on convolution or transformer architectures. Such detectors typically have a feature backbone to extract useful features from raw input images. For the remote sensing domain, a common practice among current detectors is to initialize the backbone with pre-training on ImageNet consisting of natural scenes. Fine-tuning the backbone is typically required to generate features suitable for remote-sensing images. However, this could hinder the extraction of basic visual features in long-term training, thus restricting performance improvement. To mitigate this issue, we propose a novel method named DBF (Dynamic Backbone Freezing) for feature backbone fine-tuning on remote sensing object detection. Our method aims to handle the dilemma of whether the backbone should extract low-level generic features or possess specific knowledge of the remote sensing domain, by introducing a module called ‘Freezing Scheduler’ to dynamically manage the update of backbone features during training. Extensive experiments on DOTA and DIOR-R show that our approach enables more accurate model learning while substantially reducing computational costs. Our method can be seamlessly adopted without additional effort due to its straightforward design.
Index Terms:
@@
Remote sensing object detection, feature backbone freezing, fine-tuning, transfer learning, sustainable AI.
I Introduction
With the advancement of geo-spatial and deep learning technologies, remote sensing object detection has seen astonishing progress in recent years. Modern deep learning architectures, like convolutional neural networks or transformers, have greatly improved object detection in Earth observation. These models usually include a feature backbone to extract key features for downstream tasks from the raw input images. For object detection on natural scenes, it becomes a de facto standard to initialize the backbone with pretraining on ImageNet [1] over the years, as it is one of the largest databases in visual object recognition, enabling models to learn a broad spectrum of features such as edges, textures, and shapes. Interestingly, the same research trend to use ImageNet pre-trained weights for the backbone has been observed in the remote sensing community for the following reasons. (i) Aerial and satellite imagery datasets often face challenges including limited diversity and scale due to geopolitical reasons, security concerns, and cost issues. For example, Wang et al. [2] discovered the inefficiency of solely leveraging remote sensing images for detection and segmentation tasks. (ii) Furthermore, there exists an implicit consensus that though remote sensing imagery and natural scene images have different characteristics, they may still share some common visual patterns.
Figure 1: Overview of the proposed method named DBF for efficient feature backbone fine-tuning on remote sensing object detection.
Fine-tuning the backbone with datasets for downstream remote sensing tasks is a common strategy to bridge the knowledge gap between natural and remote sensing scenes, yet it may impede overall low-level feature extraction due to the domain-restricted data, limiting performance enhancement. Hence, several researchers have concentrated on advanced remote sensing pre-training (e.g. [3, 4, 5, 6, 7, 8, 9]). For example, Fuller et al. [3] highlight the benefits of large unlabeled datasets for pre-training with masked autoencoding. Huang et al. [6] aim to learn robust representations for remote sensing understanding by simultaneously acquiring general knowledge from natural images with supervised learning and domain-specific knowledge from remote sensing images with self-supervised learning. Wang et al. [7] present multi-task pre-training for remote sensing foundation models. However, we argue that the effectiveness of such strategies is still restricted in long-term training because the weights of the backbone heavily shift further from the initial pre-trained values.
@@ I Introduction
To tackle this limitation, in this work, we introduce a novel feature backbone fine-tuning method named DBF (Dynamic Backbone Freezing) for remote sensing object detection. Inspired by recent studies of frozen settings in natural language processing (e.g. [10, 11, 12, 13]), we aim to address the dilemma of whether the backbone should extract low-level generic features or specialize in domain-specific knowledge of remote sensing. Intuitively, we implement a module named ‘Freezing Scheduler’ to dynamically control the update of the feature backbone for each epoch during training as shown in Fig. 1. The backbone receives knowledge concerning the downstream remote sensing data when the Freezing Scheduler sends a signal to open the backward route. Otherwise, the backbone becomes frozen, keeping the knowledge preserved while decreasing the computational complexity. Extensive experiments on DOTA [14] and DIOR-R [15] prove that our alternately freezing scheme can achieve better accuracy while significantly reducing GPU memory usage and training time.
@@
  • To the best of our knowledge, we first study the energy-saving feature backbone fine-tuning to address the discrepancy between upstream pre-trained weights and downstream tasks for remote sensing object detection.
  • The straightforward design of the proposed method, DBF ensures compatibility with existing deep learning frameworks, making it easy to comprehend and integrate.
  • Extensive experimental results on DOTA and DIOR-R demonstrate that our training strategy can boost remote sensing object detection in terms of prediction accuracy and computational complexity without whistles and bells.
II Methodology
In this session, we introduce a simple, but efficient feature backbone fine-tuning method namely DBF (Dynamic Backbone Fine-tuning) for remote sensing object detection.
II-A Background
@@
In computer vision, the usual practice for dense prediction tasks such as detection and segmentation had been to incorporate extra networks to avoid losing information in the feature backbone during fine-tuning (e.g. [16, 17]). After the success of the frozen setting in NLP, this concept has influenced research in computer vision tasks as well. For instance, Adapters [10] initially designed for NLP can be used in vision models to allow efficient fine-tuning without updating the entire model. Besides, Lin et al. [18] involves utilizing frozen pre-trained networks, with a task-specific head network to be fine-tuned for downstream tasks. Interestingly, Vasconcelos et al. [19] presents a fine-training scheme for detectors not to update the ImageNet-pretrained backbone. This simple strategy enables models to achieve competitive performance while significantly reducing resource usage for generic object detection on natural scenes. However, to our knowledge, only a few have focused on the frozen setting in the same context for remote sensing understanding. Motivated by the success of [19], we revisit the frozen setting for remote sensing object detection. Unlike [19], in our approach the backbone parameters are dynamically fixed or unfixed during training to meet two competing goals in remote sensing object detection: (i) generic feature extraction (for model generalization) vs. (ii) domain-specific feature extraction (for sophisticated prediction). Besides, it helps minimize the training cost as we often freeze the backbone, the most parameterized in detectors.
II-B Overview of Dynamic Backbone Freezing
@@
In this work, we propose a simple but efficient feature backbone fine-tuning method for remote sensing object detection. Our training scheme enables the model to preserve the initialized low-level generic features from natural images while harmoniously acquiring specialized knowledge in remote sensing. This is achieved by alternately freezing and unfreezing the backbone using the ‘Freezing Scheduler’ controller. The detail of DBF is presented in Fig. 2. Note that in this letter all the algorithms are written in PyTorch [20]-style pseudo code.
Figure 2: Procedure of DBF (Dynamic Backbone Freezing) in code.
@@
A deep learning-based detector is typically composed of (a) a backbone network for feature extraction, (b) a neck for feature aggregation, and (c) a detection head for predicting bounding boxes and class scores. In Fig. 2, model.backbone and model.neck correspond to (a) and (b), respectively. For (c), there are two types of heads: model.roi_head and model.bbox_head. In particular, model.roi_head is a specific type of (c) typically used in two-stage detectors like Faster R-CNN to refine the RoI (Region of Interest) proposals generated by the RPN (Region Proposal Network). Since some detectors may not have model.neck and model.roi_head, conditional statements are used to handle a runtime exception in lines 5 and 7 so that the proposed method is compatible with various architectures.
@@
Here, forward() defines the forward process of a detector model. It is the process of passing input data through the network to generate output. In this letter, we override this function for typical detector models with freezing_scheduler as shown in lines 3-4. If the condition of line 3 is true, out from model.backbone becomes detached in line 4, which means gradients for the backbone will not be computed, stopping the gradient flow to the backbone during back-propagation.
TABLE I: Experimental configuration parameters in this work.
@@
II-C Design choice of Freezing Scheduler
@@
This work assumes that the input is an epoch and the output is a 0 or 1 signal for the Freezing Scheduler. Here, 0 means allowing the backbone updates (i.e. unfreeze the backbone) while 1 represents vice versa (i.e. freeze the backbone). The flexibility of our framework enables researchers to design and implement various versions of the Freezing Scheduler. Among a wide range of variants, in this letter, we present one of the simplest forms named ‘Step Freezing Scheduler’ (Fig. 3).
@@
Figure 3: One of the simplest variants of Freezing Scheduler for DBF used in this work: ‘Step Freezing Scheduler’ in code.
@@
Intuitively, our Step Freezing Scheduler is similar to a step learning rate scheduler [23] that decreases the learning rate at specific intervals based on the users’ demand during training. In Fig. 3, $\rho$ is the only hyper-parameter to be tuned by users. Our scheduler updates the feature backbone only every $\rho$ epochs. In other words, a higher $\rho$ value implies that the backbone remains frozen more frequently, leading to increased efficiency in GPU memory usage and faster training time. Moreover, it helps to preserve the pre-trained low-level generic knowledge in the backbone while simultaneously allowing the detection model to acquire domain-specific knowledge for more precise prediction without bells and whistles.
III Experiments

### v4

# Investigating Long-term Training for Remote Sensing Object Detection

@@ Abstract
Recently, numerous methods have achieved impressive performance in remote sensing object detection, relying on convolution or transformer architectures. Such detectors typically have a feature backbone to extract useful features from raw input images. A common practice in current detectors is initializing the backbone with pre-trained weights available online. Fine-tuning the backbone is typically required to generate features suitable for remote-sensing images. While the prolonged training could lead to over-fitting, hindering the extraction of basic visual features, it can enable models to gradually extract deeper insights and richer representations from remote sensing data. Striking a balance between these competing factors is critical for achieving optimal performance. In this study, we aim to investigate the performance and characteristics of remote sensing object detection models under very long training schedules, and propose a novel method named Dynamic Backbone Freezing (DBF) for feature backbone fine-tuning on remote sensing object detection under long-term training. Our method addresses the dilemma of whether the backbone should extract low-level generic features or possess specific knowledge of the remote sensing domain, by introducing a module called ‘Freezing Scheduler’ to manage the update of backbone features during long-term training dynamically. Extensive experiments on DOTA and DIOR-R show that our approach enables more accurate model learning while substantially reducing computational costs in long-term training. Besides, it can be seamlessly adopted without additional effort due to its straightforward design. The code is available at https://github.com/unique-chan/dbf.
keywords:
@@
Remote sensing object detection, long-term training, fine-tuning, backbone freezing
1 Introduction
The rapid advancement of geospatial and deep learning technologies has revolutionized remote sensing object detection, enabling significant improvements in accuracy and efficiency. Modern object detectors rely on a feature backbone such as convolutional neural networks and transformer architectures to extract meaningful representations from raw input images [1, 2, 3]. These backbone models typically benefit from pre-trained weights on large-scale datasets such as ImageNet [4], which serve as a foundation for learning low-level features like edges and textures. Fine-tuning these pre-trained backbones with remote sensing datasets is a common practice to adapt to the unique characteristics of aerial and satellite imagery.
However, the process of fine-tuning under long-term training schedules introduces unique challenges and opportunities for remote sensing object detection. On the one hand, prolonged training can lead to overfitting, especially in feature backbones, hindering the extraction of fundamental visual features critical for generalization. On the other hand, such extended schedules provide a valuable opportunity to acquire domain-specific and enriched knowledge from data, potentially enhancing the model’s ability to detect intricate patterns unique to remote sensing imagery. For example, Beyer et al. [5] empirically found that prolonged training with appropriate data augmentation strategies may be important to achieve better accuracy. This trade-off between preserving generic features and specializing for the domain presents a critical challenge in designing effective training strategies for feature backbones.
In addition, many modern object detection models are released with pre-trained weights based on natural image datasets such as ImageNet. To adapt these models to remote sensing tasks, fine-tuning is essential; however, as the model size increases, this process becomes increasingly resource-intensive. In particular, training large backbones from scratch or even fully fine-tuning them requires substantial GPU memory and computation time, which can hinder practical deployment in remote sensing scenarios [6, 7, 8].
@@ 1 Introduction
To tackle this limitation, we introduce a novel feature backbone fine-tuning method named DBF (Dynamic Backbone Freezing) for remote sensing object detection under long-term training. Our primary objective is to examine the performance and behavioral traits of remote sensing object detection models when subjected to prolonged training schedules with our fine-tuning approach. Inspired by recent studies of fine-tuning with frozen settings in natural language processing (e.g. [9, 10, 11, 12]), we aim to address the dilemma of whether the backbone should extract low-level generic features or specialize in domain-specific knowledge of remote sensing. Intuitively, we implement a module named ‘Freezing Scheduler’ to dynamically control the update of the feature backbone for each epoch during training as shown in Fig. 1. The backbone receives knowledge concerning the downstream remote sensing data when the Freezing Scheduler sends a signal to open the backward route. Otherwise, the backbone becomes frozen, keeping the knowledge preserved while decreasing the computational complexity. Extensive experiments on DOTA [13] and DIOR-R [14] prove that our alternately freezing scheme can achieve better accuracy while significantly reducing GPU memory usage and training time, as indicated in Fig. 2 and Table 2.
Figure 1: Illustration of training remote sensing object detectors with the proposed framework named Dynamic Backbone Freezing (DBF). DBF allows the backbone to receive knowledge from the downstream data when Freezing Scheduler sends a signal to open the backward route. This ensures robust prediction by preventing over-fitting while also significantly saving training costs.
Figure 2: Impact of the proposed method on remote sensing object detection. The above visualization indicates the performance of DOTA and DIOR-R benchmark datasets using models using FCOS with ResNet-50. Under long-term training, our method achieves the highest prediction accuracy (mAP) while significantly reducing computational costs (FLOPs), compared to existing methods. Besides FCOS, Faster R-CNN and RetinaNet also show similar trends.
@@
  • We present DBF (Dynamic Backbone Freezing), a novel fine-tuning strategy that enables efficient adaptation of pre-trained models to the remote sensing domain under long-term training scenarios.
  • DBF significantly reduces training cost—such as GPU memory usage and FLOPs—while preserving or even enhancing detection performance, especially for large-scale models for remote sensing object detection.
  • The design of DBF is simple and framework-agnostic, ensuring seamless integration into existing remote sensing object detection models.
  • We conduct extensive experiments on two well-known benchmarks for remote sensing object detection, including DOTA and DIOR-R, demonstrating that DBF achieves consistent gains over both full training and frozen-backbone baselines.
The remainder of this paper is organized as follows: Section 2 reviews related work to provide context and highlight the contributions of this study. Section 3 introduces the proposed methodology, including the design and implementation of DBF. Section 4 and 5 present experimental results and a detailed analysis of the method’s performance. Finally, Section 6 concludes the study and outlines potential directions for future research.
2 Related work
2.1 Pre-training and fine-tuning for remote sensing object detection
Unlike natural scene images in datasets such as ImageNet, remote sensing images often exhibit distinct domain characteristics, including small object sizes, low contrast, and degradation caused by atmospheric or sensor-related conditions such as haze, fog, and rain. These factors substantially hinder detection accuracy, making image enhancement and restoration critical pre-processing steps.
Nevertheless, even with enhanced input quality, domain-specific fine-tuning remains essential, as feature backbones pre-trained on natural image datasets still require adaptation to effectively capture the unique spatial and spectral characteristics of remote sensing scenes.
Modern deep learning-based object detectors usually include a feature backbone to extract key features for downstream tasks from the raw input images. For object detection on natural scenes, it becomes a de facto standard to initialize the backbone with pretraining on ImageNet [4] over the years, as it is one of the largest databases in visual object recognition, enabling models to learn a broad spectrum of features such as edges, textures, and shapes. Interestingly, the same research trend to use ImageNet pre-trained weights for the backbone has been observed in the remote sensing community for the following reasons. (i) Aerial and satellite imagery datasets often face challenges including limited diversity and scale due to geopolitical reasons, security concerns, and cost issues. For example, Wang et al. [15] discovered the inefficiency of solely leveraging remote sensing images for detection and segmentation tasks. (ii) There exists an implicit consensus that though remote sensing imagery and natural scene images have different characteristics, they may still share some common visual patterns. (iii) ImageNet pre-trained weights are readily accessible through various deep learning frameworks online available.
Fine-tuning the backbone with datasets for downstream remote sensing tasks is a common strategy to bridge the knowledge gap between natural and remote sensing scenes, yet it may impede overall low-level feature extraction due to the domain-restricted remote sensing data, limiting performance enhancement. Hence, several researchers have concentrated on advanced remote sensing pre-training (e.g. [16, 17, 18, 19, 20, 21, 22]).
For instance, the benefits of large unlabeled datasets for pre-training with masked autoencoding have been demonstrated [16]. Robust representations for remote sensing understanding can be learned by simultaneously acquiring general knowledge from natural images via supervised learning and domain-specific knowledge from remote sensing images through self-supervised learning [19]. Multi-task pre-training has also been introduced to construct remote sensing foundation models [20]. Semi-supervised object detection using uncurated data has been proposed [23]. A transformation-invariant few-shot detection model improves generalization in limited data regimes [24]. Point-level supervision significantly reduces annotation cost while maintaining segmentation accuracy [25].
However, we argue that the effectiveness of such pretraining strategies is still restricted in long-term training because the weights of the backbone heavily shift further from the initial pre-trained values. To mitigate this issue, we focus on designing a feature backbone fine-tuning method for remote sensing object detection under long-term training schemes.
2.2 Freezing backbones in deep learning
@@
In addition, recent efforts have focused on enhancing detection performance while reducing computational complexity by eliminating attention mechanisms. For example, Gao et al. [26] propose an attention-free global multiscale fusion network tailored for remote sensing object detection, demonstrating efficient feature extraction without the computational overhead of self-attention.
Beyond parameter-efficient tuning strategies, some researchers have explored alternative architectures inspired by neuromorphic computing. For instance, Li et al. [27] present a spiking neural network-based object detector, significantly reducing energy consumption while maintaining detection performance. Such approaches align with the increasing demand for lightweight and efficient models in resource-constrained remote sensing environments.
@@
In computer vision, the usual practice for dense prediction tasks such as detection and segmentation had been to incorporate extra networks to avoid losing information in the feature backbone during fine-tuning (e.g. [28, 29]). After the success of the frozen setting in NLP, this concept has influenced research in computer vision tasks as well. Lin et al. [30] involves utilizing frozen pre-trained networks, with a task-specific head network to be fine-tuned for downstream tasks. Interestingly, Vasconcelos et al. [31] presents a fine-training scheme for detectors not to update the ImageNet-pretrained backbone. This simple strategy enables models to achieve competitive performance while significantly reducing resource usage for generic object detection on natural scenes. However, to our knowledge, only a few have focused on the frozen setting in the same context for remote sensing understanding. Therefore, in this paper, we revisit the frozen setting for remote sensing object detection.
2.3 Recent object detection paradigm for remote sensing
Oriented object detection refers to a specialized task where the goal is to locate and classify objects in an image while accounting for their orientation (rotation). Unlike traditional object detection, which uses AABBs (Axis-Aligned Bounding Boxes) to enclose objects, oriented object detection employs OBBs (Oriented Bounding Boxes) that align with the object’s true orientation in the image [32].
Traditionally, convolution neural network-based oriented object detectors can be broadly classified into two categories: one-stage methods and two-stage methods. One-stage object detection methods frame object detection as a single regression problem. These models directly predict class probabilities and bounding box coordinates from feature maps, bypassing the need for a region proposal stage. By optimizing detection end-to-end, one-stage detectors achieve significantly faster inference speeds, making them highly suitable for real-time applications. Traditionally, these methods employ anchor-based mechanisms, where predefined anchor boxes of various scales and aspect ratios guide the network to predict bounding boxes. While efficient, anchor-based one-stage detectors can suffer from challenges such as anchor design sensitivity. To address this, anchor-free one-stage detectors, such as FCOS [33], Oriented Rep [34], and CFA [35], propose bounding boxes by directly predicting object center points or key points, eliminating the dependency on predefined anchors. Two-stage object detection methods, typified by the R-CNN family (e.g., Faster R-CNN [36], Oriented R-CNN [37] and ReDet [38]), follow a decoupled pipeline. The first stage generates candidate region proposals (via an RPN in anchor-based methods), while the second stage refines these proposals and predicts class labels. Anchor-based two-stage detectors have long been the standard, with their region proposal networks relying on anchors to localize potential objects effectively.
Recently, transformer-based methods have emerged as a groundbreaking approach in object detection, shifting the paradigm from traditional convolutional designs to sequence modeling. These models leverage self-attention mechanisms and end-to-end optimization, as seen in DETR (DEtection TRansformer) [39] and Deformable DETR [40]. CLT-DETR [41] is the first adaptation of DETR for oriented object detection. AO2-DETR [42] enhances Deformable DETR by introducing an oriented box generation and refinement module, improving localization accuracy. D2Q-DETR [43] further refines this paradigm with dynamic queries that reduce object tokens progressively, balancing accuracy and efficiency. OrientedFormer [3] introduces Gaussian positional encoding, Wasserstein self-attention, and oriented cross-attention to detect arbitrarily oriented objects in remote sensing images.
Building upon these advances, recent studies have explored efficient backbone architectures tailored to ultra-high-resolution remote sensing imagery. DynamicVis [44] proposes a selective state-space backbone that adaptively attends to informative regions while suppressing less relevant areas. This design achieves high performance with lower complexity than traditional transformer-based models, offering a promising direction for scalable and generalizable remote sensing backbones.
In addition, visual foundation models have been increasingly adopted to adapt large-scale pre-trained vision backbones to remote sensing tasks. Rather than full model retraining, these approaches often utilize efficient fine-tuning strategies. For example, RSPrompter [45] builds on the Segment Anything Model (SAM), incorporating LoRA-based prompt learning for parameter-efficient instance segmentation. Similarly, RSRefSeg2 [46] introduces a CLIP-SAM integrated framework with dual-stage referring prompts for open-vocabulary segmentation. These methods demonstrate that with appropriate prompt design and cross-modal alignment, visual foundation models can be flexibly and efficiently adapted to a wide range of remote sensing applications.
3 Proposed approach
In this section, we introduce a simple, but efficient feature backbone fine-tuning method namely Dynamic Backbone Fine-tuning (DBF) for remote sensing object detection, particularly in long-term training.
3.1 Motivation
Compared to the existing freezing methods like [31], which adopt a static freezing strategy, our approach introduces dynamic adjustments to the trainability of backbone parameters during training. This innovation enables the model to effectively balance two competing objectives in remote sensing object detection: (i) generic feature extraction, which ensures robust generalization across diverse input types (- by unfreezing -), and (ii) domain-specific feature extraction, which facilitates accurate and specialized predictions (- by freezing -). Previous approaches, such as [31], often adopt a static freezing mechanism where the backbone remains either completely trainable or entirely frozen throughout the training process. While this reduces computational overhead, it may fail to effectively balance generalization and specialization. Fully frozen backbones risk underfitting to domain-specific characteristics, while fully trainable backbones may lose essential low-level generic features from pre-trained weights.
In this study, ‘long-term’ training refers to schedules extending beyond 200 epochs, allowing models to gradually acquire enriched domain-specific knowledge while preserving generic features through DBF. In existing papers, the models are usually trained for 12 epochs with remote sensing object detection benchmarks such as DOTA and DIOR-R [47].
3.2 Overview of Dynamic Backbone Freezing
@@
In this work, we propose a simple but efficient feature backbone fine-tuning method for remote sensing object detection. Our training scheme enables the model to preserve the initialized low-level generic features from natural images while harmoniously acquiring specialized knowledge in remote sensing. This is achieved by alternately freezing and unfreezing the backbone using the ‘Freezing Scheduler’ controller.
Figure 3: Visual illustration of long-term training strategies for remote sensing object detection over 400 epochs: (a) Full Training, (b) Frozen Backbone, (c) Ours (DBF).
More precisely, DBF inherently aligns with the dual objectives by:
  • Freezing the backbone to prioritize generalization and prevent overfitting to limited domain-specific data.
  • Unfreezing the backbone intermittently to allow fine-tuning on specialized features as the training progresses and domain-specific patterns become more apparent.
The detail of DBF is presented in Algorithm 1. Note that in this paper all the algorithms are written in PyTorch [48] style.
1def forward(input):
2 out = model.backbone(input)
3 if freezing_scheduler(epoch): # proposed
4 out = out.detach()
5 if model.neck:
6 out = model.neck(out)
7 if model.roi_head:
8 out = model.roi_head(out)
9 out = model.bbox_head(out)
10 return out
ALGORITHM 1: Training with Dynamic Backbone Freezing.
@@
A deep learning-based detector is typically composed of (a) a backbone network for feature extraction, (b) a neck for feature aggregation, and (c) a detection head for predicting bounding boxes and class scores. In Algorithm 1, backbone and neck correspond to (a) and (b), respectively. For (c), there are two types of heads: roi_head and bbox_head. In particular, roi_head is a specific type of (c) typically used in two-stage detectors like Faster R-CNN to refine the RoI (Region of Interest) proposals generated by the RPN (Region Proposal Network). Since some detectors may not have neck and roi_head, conditional statements are used to handle a runtime exception in lines 5 and 7 so that the proposed method is compatible with various architectures.
@@
Here, forward() defines the forward process of a detector model. It is the process of passing input data through the network to generate output. In this paper, we override this function for typical detector models with freezing_scheduler as shown in lines 3-4. If the condition of line 3 is true, out from backbone becomes detached in line 4, which means gradients for the backbone will not be computed, stopping the gradient flow to the backbone during back-propagation.
Table 1: Experimental configuration parameters in this work.
@@
3.3 Design choice of Freezing Scheduler
@@
This work assumes that the input to the Freezing Scheduler is the current epoch, and the output is a binary signal indicating whether the backbone should be frozen. Specifically, a signal of 0 allows backbone updates (i.e., unfreezes the backbone), while a signal of 1 freezes it. Using this scheduler, we can dynamically control the training process by alternating between freezing and unfreezing the backbone, as illustrated in Fig. 3. The flexibility of our framework enables researchers to design and implement various versions of the Freezing Scheduler according to their needs. Among these possible variants, we propose a simple yet effective approach called the “Step Freezing Scheduler,” which is described in Algorithm 2.
1def freezing_scheduler(epoch):
2 # ρ: an integer greater than 1 and
3 # less than total epochs
4 if epoch % ρ == 0:
5 return 0 # unfreezing
6 return 1 # freezing
@@
ALGORITHM 2: One of the simplest variants of Freezing Scheduler for DBF in this work: ‘Step Freezing Scheduler’.
@@
Intuitively, our Step Freezing Scheduler is similar to a step learning rate scheduler [51] that decreases the learning rate at specific intervals based on the users’ demand during training. In Algorithm 2, $\rho$ is the only hyper-parameter to be tuned by users. Our scheduler updates the feature backbone only every $\rho$ epochs. In other words, a higher $\rho$ value implies that the backbone remains frozen more frequently, leading to increased efficiency in GPU memory usage and faster training time. Moreover, it helps to preserve the pre-trained low-level generic knowledge in the backbone while simultaneously allowing the detection model to acquire domain-specific knowledge for more precise prediction.
4 Experiments